OpenStudy (anonymous):

Given that f(x) = 3x + 1 and g(x) = the quantity of 4x plus 2 divided by 3, solve for g(f(0)). 2 6 8 9

OpenStudy (anonymous):

@UnkleRhaukus can you help

OpenStudy (unklerhaukus):

can you turn the expression for g(x) into mathematical symbols g(x) = the quantity of 4x plus 2 divided by 3 =

OpenStudy (anonymous):

g(x)=4x+2/3

OpenStudy (unklerhaukus):

' the quantity ' bit means you need some parenthesis g(x)=(4x+2)/3

OpenStudy (unklerhaukus):

The question is looking for g(f(0)) lets find f(0) first; f(x) = 3x + 1 subbing in zero for x, we get f(0) = 3(0) + 1 = what does this simplify to ?

OpenStudy (anonymous):

f(0)=4

OpenStudy (anonymous):

no wait not 4

OpenStudy (anonymous):

1

OpenStudy (unklerhaukus):

careful that 3(0) really mean 3 x 0

OpenStudy (unklerhaukus):

yeah f(0) = 1

OpenStudy (unklerhaukus):

so all we have left to do is find g(f(0)) = g(1) and we already have g(x) = (4x+2)/3 so sub in x=1 g(1) =

OpenStudy (anonymous):

g(1)=(4(1)+2)/3 =(4+2)/3 = 6/3 =2 g(1)=2

OpenStudy (unklerhaukus):

\[\huge\color{red}\checkmark\]

OpenStudy (anonymous):

thank you

OpenStudy (unklerhaukus):

good work